# Foundation Models in Electrocardiogram: A Review

### v2

@@ Abstract
The electrocardiogram (ECG) is ubiquitous across various healthcare domains, such as cardiac arrhythmia detection and sleep monitoring, making ECG analysis critically essential. Traditional deep learning models for ECG are task-specific, with a narrow scope of functionality and limited generalization capabilities. Recently, foundation models (FMs), also known as large pre-training models, have fundamentally reshaped the scheme of model design and representation learning, enhancing the performance across a variety of downstream tasks. This success has drawn interest in the exploration of FMs to address ECG-based medical challenges concurrently. This survey provides a timely, comprehensive and up-to-date overview of FMs for large-scale ECG-FMs. First, we offer a brief background introduction to FMs. Then, we discuss the model architectures, pre-training methods, and adaptation approaches of ECG-FMs from a methodology perspective. Despite the promising opportunities of ECG-FMs, we also outline the challenges and potential future directions. Overall, this survey aims to provide researchers and practitioners with insights into the research of ECG-FMs on theoretical underpinnings, domain-specific applications, and avenues for future exploration.
1 Introduction
@@
Foundation models (FMs), representing a new generation of AI models trained on massive and diverse datasets, have grown rapidly across a range of fields, including language understanding [51, 85], computer vision [120], and multimodal tasks [26, 128]. Notable examples include OpenAI’s GPT-4 [2], a large language model that demonstrates state-of-the-art capabilities in natural language processing tasks like text generation and summarization, and Meta’s LLaMA-series [117, 22], which pushes the boundaries of open-source large-scale models by offering high-performance models trained on diverse data. These models are designed to generalize across multiple tasks, showing the potential for flexibility, adaptability, and scalability in a wide variety of applications [37, 83].
In the medical domain, and particularly in electrocardiography (ECG), FMs have also garnered significant attention [84, 89]. Traditionally, ECG models were built with task-specific objectives, such as detecting a specific cardiovascular condition like atrial fibrillation (AF) or premature ventricular contractions (PVCs) [89, 36]. These early models required manually labeled datasets for each condition and were limited by their narrow scope.
In contrast, ECG foundation models (ECG-FMs) present a more versatile solution, offering several advantages [104]. On the one hand, they can perform multiple tasks, such as arrhythmia detection, HRV analysis, and even more complex tasks like predicting heart failure or stratifying patients based on risk profiles [126]. For example, FMs trained on large ECG datasets, like the PTB-XL [111] or PhysioNet’s MIMIC-III [47] waveform database, have been shown to generalize across multiple tasks with high accuracy. These models leverage the diversity of the data, learning features that are transferable across tasks and patient populations, making them better suited for real-world clinical applications where data variability is common [94]. On the other hand, ECG foundation models can incorporate multi-modal data, integrating ECG signals with other physiological or clinical data such as photoplethysmography (PPG) [58] or electronic health records (EHRs) [16], to provide a more comprehensive picture of patient health [3].
Despite the promising prospects and rapid development of ECG-FMs, a systematic analysis of ECG-FMs from a methodological standpoint has been notably absent in prior literature. To bridge this gap, this survey aims to provide a comprehensive methodological analysis of foundation models for ECG-based clinical diagnosis. This examination will center on scrutinizing their architecture design, pre-training paradigms, and adaptation techniques. Through this endeavor, we seek to illuminate an overall picture of core elements in ECG-FMs, thereby enhancing comprehension regarding the rationale behind their efficacy and the mechanisms driving their substantial potential in ECG analysis.
In contrast to previous surveys which concentrated on general-purpose FMs or FMs covering a wide scope of medical fields, this article focuses on the specialized domain of ECG analysis. In this article, we adopt the settings in  [65] to categorize the ECG-FMs into Transformer-based models and non-Transformer-based models. Notably, while more than half of the existing studies rely on simpler deep learning models (e.g., CNN and ResNet) rather than large-scale models (e.g., BERT and LLaMA) that have millions of parameters, these simpler models hold significant potential for further development. By increasing the scale of data and enhancing model depth, they can evolve into large frameworks. Therefore, we regard all of these models as foundation models in this article, underscoring their relevance and versatility within the field. We summarize the developmental roadmap of current ECG-FMs in Figure 1, to inspire advancements and deepen comprehension in the continuously evolving landscape of FMs for ECG analysis. In brief, our major contributions are summarized as follows.
@@
  • Comprehensive and up-to-date review. We thoroughly review the existing literature on ECG-FMs, and provide a comprehensive and up-to-date review on FMs for ECG analysis, encompassing ECG-related datasets and benchmarks, along with a wide spectrum of tasks for ECG analysis.
  • Novel methodology-centric taxonomy. We present a novel taxonomy that delivers a thorough analysis from a methodological standpoint of FMs for ECG analysis, accompanied by a detailed depiction of the general architectures and pipeline design.
  • Current Challenges and future research opportunities. We discuss the most challenging problems in ECG-FM research from both methodological and applicative perspectives. Additionally, we explore and emphasize potential future directions for enhancing ECG analysis through the applications of foundation models, encouraging practitioners to undertake further investigations in this field.
@@
The rest of the article is organized as follows. Section 2 illustrates the search strategy. Section 3 describes the brief background of FMs. Section 4 presents the methodology of ECG-FMs. Section 5 discusses the existing challenges and promising directions of ECG-FMs, and Section 6 concludes the article.
@@
2 Search Strategy
All the candidate articles were collected by a comprehensive search of three databases, including PubMed, arXiv, and Google Scholar using the query conditions: ‘ECG foundation model’ $||$ ‘foundation model in ECG analysis’$||$ ‘ECG self-supervised learning’. After removing duplicated papers, we included all research papers in English. Since this review mainly focuses on methodology, we included 46 papers describing methods for foundation models in ECG analysis using deep learning algorithms.
3 Background of Foundation Models
FMs are large-scale deep-learning models trained on diverse and extensive datasets in a self-supervised or semi-supervised manner [12]. FMs generally follow the pre-training and fine-tuning paradigm, which enables the model to learn a broad range of knowledge and patterns and to leverage such prior information in various downstream tasks and applications. FMs only require considerably small-scale custom data in the fine-tuning phase, showcasing remarkable adaptability and efficiency [125]. FMs have been adopted successfully across domains such as CV, NLP, and healthcare. As the data or model size massively scales up and the available data modality dramatically increases, FMs have evolved from the initial pre-trained models like BERT to contemporary large language models (LLMs), vision foundation models (VFMs), vision-language models (VLMs), and time-series foundation models (TSFMs).
Large Language Models.
In the sphere of NLP, LLMs have fundamentally reshaped our understanding of syntactic and semantic information extracted from input text. Autoregressive-based LLMs, exemplified by powerful models like GPT-3 [81], PaLM [15], and ChatGPT [14, 110], have redefined text generation and question-answering capabilities by predicting subsequent words in a sequence and producing contextually coherent outputs. Additionally, the emergence of the groundbreaking LLM, LLaMA [117], represents a significant advancement, striking a balance between high performance, efficiency, and accessibility through innovative encoding tokenization and memory-efficient attention techniques. On the other hand, autoencoder-based LLMs, such as BERT [19] and its remarkable variants like T5 [99], RoBERTa [76], DistilBERT [102], and ELECTRA [17], have redefined the analysis of contextual relationships within text by leveraging bidirectional context consideration. This makes them an indispensable asset for a wide range of NLP tasks, from sentiment analysis to machine translation.
Vision Foundation Models.
@@ 3 Background of Foundation Models
Vision Language Models.
@@
Time Series Foundation Models.
The increasing richness of time-series data and its wide range of applications have thrust TSFMs into the spotlight as a promising direction for time series analysis [46, 65]. TSFMs aim to harness the power of the FM paradigm to develop generalized models adept at understanding and forecasting time series data across diverse domains. By leveraging large-scale time series datasets, TSFMs hold the promise of attaining superior performance across a spectrum of time series tasks, offering a unified framework that can expedite research and application development in this field. Based on the nature of time series, TSFMs can be classified into general-purpose TSFMs dealing with standard time series and domain-specific TSFMs focusing on particular tasks such as traffic forecasting and clinical diagnosis. General-purpose TSFMs predominantly utilize single or multiple data modalities to craft robust models targeting specific time series tasks, such as Voice2Series [129], TS2Vec [137], and SimMTM [20] for classification; UniTime [74] for cross-domain prediction; UniTS [29] for multi-task learning; and TimesFM [18], TimeGPT-1 [30], TTMs [24], and TSMixer [23] for zero-shot forecasting; Domain-specific TSFMs have distinct focuses and architectures tailored to the requirements of the downstream tasks. In the transportation sector, TSFMs typically employ graph structures, such as GNNs, to capture the spatio-temporal interactions within dynamic traffic systems. For instance, STGCL [75] and SPGCL [64] explore the integration of contrastive learning into spatio-temporal graph forecasting, while ST-LLM [68] combines spatio-temporal information with a partially frozen LLM to enhance traffic predictions. Additionally, TSFMs in the medical domain usually place greater emphasis on capturing the hierarchical characteristics of patient data. For example, FORMED [45] can easily capture domain knowledge and be adapted to unseen cohorts with minimum training by repurposing a pre-trained TimesFM. COMET [123] utilizes a multi-granularity contrastive learning framework to exploit data consistencies across different levels (e.g., observation-level and patient-level) in various medical time series, showcasing its potential benefits in complex disease diagnosis.
4 Methodology Perspective
In this section, we introduce the FMs for ECG analysis from a methodology perspective, focusing on architecture design (including backbone model and Transformer structure), and pipeline (including pre-training paradigm and fine-tuning paradigm). This discussion aims to shed light on the mechanisms of the FMs driving the effectiveness and adaptability on different ECG-based tasks.

### v3

# A Systematic Review on Foundation Models for Electrocardiogram Analysis: Initial Strides and Expansive Horizons

@@ Abstract
Electrocardiogram (ECG) is widely used in healthcare applications, such as arrhythmia detection and sleep monitoring, making accurate ECG analysis critically essential. Traditional deep learning models for ECG are task-specific, with limited generalization and narrow functionality. Foundation models (FMs), or large pre-training models, have recently advanced representation learning, enabling strong performance across diverse tasks and motivating their adoption for ECG analysis. Here, we present the first comprehensive review dedicated to ECG foundation models (ECG-FMs). We map the current landscape of architectures, pre-training paradigms, and adaptation strategies, and critically examine their strengths, limitations, and clinical potential. By consolidating this emerging field, we aim to accelerate the development of robust, generalizable ECG-FMs and chart future directions for their integration into healthcare practice.
Introduction
@@
Foundation models (FMs), a new generation of AI models trained on massive and diverse datasets, have advanced rapidly across multiple domains, including language understanding [1, 2], computer vision (CV) [3], and multimodal tasks [4, 5]. Notable examples include OpenAI’s GPT-4 [6], a large language model that demonstrates state-of-the-art performance in natural language processing (NLP) tasks such as text generation and summarization, and Meta’s LLaMA series [7, 8], which extends the frontier of open-source large-scale modeling by offering high-performance models trained on diverse data. These models are designed to generalize across tasks, highlighting their flexibility, adaptability, and scalability across a wide range of applications [9, 10].
In recent years, FMs have attracted growing interest in the medical domain, including electrocardiogram (ECG) analysis [11, 12]. Conventional ECG models are typically designed for narrow, task-specific objectives, relying heavily on labeled datasets and constrained by limited generalizability. By contrast, ECG foundation models (ECG-FMs) offer a more flexible and broadly applicable framework with several notable advantages [13]. First, ECG-FMs support a wide range of tasks, from general cardiac disease diagnosis, such as arrhythmia detection [14] and heart rate variability (HRV) analysis [15], to grounded ECG understanding, including predicting heart failure [16] and stratifying patient risk profiles [17]. Through pre-training on large and heterogeneous ECG datasets, these models learn transferable representations that exhibit robustness across diverse patient populations, thereby enhancing their utility in real-world clinical environments characterized by substantial variability in data distributions [18]. Second, ECG-FMs enable multimodal integration by combining ECG signals with complementary data sources such as photoplethysmography (PPG)[19] or electronic health records (EHRs)[20]. This integration facilitates a more holistic assessment of patient health and provides richer diagnostic insights [21].
Despite these advances, prior literature lacks a systematic methodological survey of ECG-FMs. To address this gap, this article provides a comprehensive review that analyzes ECG-FMs from three perspectives: architecture design, pre-training paradigms, and adaptation techniques. This analysis seeks to clarify the mechanisms underpinning the effectiveness of ECG-FMs and to highlight their potential in ECG analysis.
Unlike earlier surveys that focus on general-purpose FMs [22] or broadly cover medical foundation models [11, 23, 24], our work concentrates specifically on ECG analysis. Following the categorization in the survey [25], we devide ECG-FMs into Transformer-based models and non-Transformer-based models. Notably, while more than one third of existing studies employ relatively simple architectures such as convolutional neural networks (CNNs) or ResNets, rather than large-scale models like BERTs or LLaMAs, these models still hold significant potential for expansion. With larger datasets and deeper architectures, they can evolve into robust FM frameworks. In this article, we therefore consider both large-scale and simple deep learning models as part of the ECG-FM landscape, underscoring their relevance and versatility. To inspire further innovation, we summarize the development roadmap of current ECG-FMs in Figure 1. Our contributions can be summarized as follows:
@@
  • Comprehensive and up-to-date review. We provide a thorough and current review of ECG-FMs, including datasets, benchmarks, and a broad spectrum of tasks for ECG analysis.
  • Novel methodology-centric taxonomy. We introduce a novel taxonomy that examines ECG-FMs from a methodological perspective, with a detailed discussion of architectural design and pipeline workflows.
  • Current challenges and future research opportunities. We identify key challenges in ECG-FM research from both methodological and clinical viewpoints and outline promising directions for future work, encouraging further exploration of foundation models for ECG analysis.
@@
The rest of the article is organized as follows. Section 2 outlines the literature search strategy. Section 3 introduces the background of FMs. Section 4 reviews the architectures of ECG-FMs, and Section 5 introduces their general development pipeline. Section 6 discusses existing challenges and highlights promising research directions. Finally, Section 7 concludes the article.
@@
Literature search strategy
We conduct a systematic review of foundation models for ECG analysis following the Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) guidelines [26]. Our review focused on English articles published in recent five years, sourced from databases such as PubMed, arXiv, Google Scholar, and publisher platforms like Elsevier, ACM, and IEEE Xplore. To refine our search, we used the following search query: (“foundation model" OR “large pre-training model" OR “deep learning" OR “self-supervised learning") AND (“ECG" OR “electrocardiogram" OR “ECG analysis"). Our search yielded a total of 3330 studies, and 66 were retrieved for full-text review after duplicates removal and title and abstract screening. In total, 60 were selected for inclusion in the systematic review, as shown in Figure 2. Studies offering differing perspectives or contradictory evidence were carefully selected to ensure a balanced and comprehensive view of ECG-FMs. The summary of the included studies are shown in Table 1.
Figure 2: PRISMA diagram of the literature review process. We retrieved 60 publications among the 3330 papers that are collected from the six academic platforms.
Background of foundation models
FMs are large-scale deep-learning models trained on diverse and extensive datasets in a self-supervised or semi-supervised manner [27]. FMs generally follow the pre-training and fine-tuning paradigm, which enables the model to learn a broad range of knowledge and patterns and to leverage such prior information in various downstream tasks and applications. FMs only require considerably small-scale custom data in the fine-tuning phase, showcasing remarkable adaptability and efficiency [28]. FMs have been adopted successfully across domains such as CV, NLP, and healthcare. As the data or model size massively scales up and the available data modality dramatically increases, FMs have evolved from the initial pre-trained models like BERT to contemporary language foundation models (LFMs), vision foundation models (VFMs), vision-language foundation models (VLFMs), and time-series foundation models (TSFMs).
Language foundation models
In the sphere of NLP, large language models (LLMs) have fundamentally reshaped our understanding of syntactic and semantic information extracted from input text. Autoregressive-based LLMs, exemplified by powerful models like GPT-4 [6], Qwen2.5 [29], and DeepSeek-v3 [30], have redefined text generation and question-answering capabilities by predicting subsequent words in a sequence and producing contextually coherent outputs. Additionally, the emergence of the groundbreaking LLM, LLaMA3 [8], represents a significant advancement, striking a balance between high performance, efficiency, and accessibility through innovative encoding tokenization and memory-efficient attention techniques. On the other hand, autoencoder-based LLMs, such as BERT [31] and its remarkable variants like T5 [32], RoBERTa [33], DistilBERT [34], and ELECTRA [35], have redefined the analysis of contextual relationships within text by leveraging bidirectional context consideration. This makes them an indispensable asset for a wide range of NLP tasks, from sentiment analysis to machine translation.
Vision foundation models
@@
Vision-language foundation models
@@
Time series foundation models
The increasing richness of time-series data and its wide range of applications have thrust TSFMs into the spotlight as a promising direction for time series analysis [48, 25]. TSFMs aim to harness the power of the FM paradigm to develop generalized models adept at understanding and forecasting time series data across diverse domains. By leveraging large-scale time series datasets, TSFMs hold the promise of attaining superior performance across a spectrum of time series tasks, offering a unified framework that can expedite research and application development in this field. Based on the nature of time series, TSFMs can be classified into general-purpose TSFMs dealing with standard time series and domain-specific TSFMs focusing on particular tasks such as traffic forecasting and clinical diagnosis. General-purpose TSFMs predominantly utilize single or multiple data modalities to craft robust models targeting specific time series tasks, such as Voice2Series [49], TS2Vec [50], and SimMTM [51] for classification; UniTime [52] for cross-domain prediction; UniTS [53] for multi-task learning; and TimesFM [54], TimeGPT-1 [55], TTMs [56], and TSMixer [57] for zero-shot forecasting. Domain-specific TSFMs have distinct focuses and architectures tailored to the requirements of the downstream tasks. In the transportation sector, TSFMs typically employ graph structures, such as graph neural networks (GNNs), to capture the spatio-temporal interactions within dynamic traffic systems. For instance, STGCL [58] and SPGCL [59] explore the integration of contrastive learning into spatio-temporal graph forecasting, while ST-LLM [60] combines spatio-temporal information with a partially frozen LLM to enhance traffic predictions. Additionally, TSFMs in the medical domain usually place greater emphasis on capturing the hierarchical characteristics of patient data. For example, FORMED [61] can easily capture domain knowledge and be adapted to unseen cohorts with minimum training by repurposing a pre-trained TimesFM. COMET [62] utilizes a multi-granularity contrastive learning framework to exploit data consistencies across different levels (e.g., observation-level and patient-level) in various medical time series, showcasing its potential benefits in complex disease diagnosis.
Architecture design of ECG-FMs
As shown in Figure 3, we investigate the architectures of existing ECG-FMs, focusing on their backbone models and Transformer modes. This section aims to elucidate the underlying mechanisms that shape the capabilities of ECG-FMs and their applications across diverse ECG-based scenarios.